# MUSCLE: A Model Update Strategy for Compatible LLM Evolution

### v1

@@ 4 Knowledge Transfer
[…]
@@
Most knowledge distillation works consider the distillation from a trained teacher to an untrained student Tian et al. (2022); Rajasegaran et al. (2020). Recent work (Roth et al., 2024) tackles the goal of knowledge transfer between pretrained student and teacher models while retaining student knowledge gained a priori, and show that standard knowledge distillation between pretrained models struggles to transfer knowledge without performance drops for many teacher-student pairs. Complementary to this work that focuses on performance and maintaining prior knowledge, we tackle compatibility to previous models through knowledge transfer.
@@ 4.1 Model Update Strategy for Compatible LLM Evolution (MUSCLE)
When the base model is updated, we aim to train a task-specific fine-tuned model, $\mathcal{M}^{C}_{v2}$, that has the accuracy benefits of $\mathcal{M}_{v2}$, but with most compatibility with $\mathcal{M}_{v1}$. We obtain $\mathcal{M}^{C}_{v2}$ by training a compatibility adapter applied to the base model $\mathcal{M}^{\text{base}}_{v2}$. We use knowledge from task-specific fine-tuned models $\mathcal{M}_{v1}$ and $\mathcal{M}_{v2}$ when training $\mathcal{M}^{C}_{v2}$. $\mathcal{M}_{v2}$ typically has increased prediction capabilities over $\mathcal{M}_{v1}$ (due to improvements in the base model), but $\mathcal{M}_{v1}$ has information on already correctly predicted tokens or instances that we want to minimize regression towards.
We initialize the compatibility adapter with the task-specific adapter of $\mathcal{M}_{v2}$, and further fine-tune it (using the task training dataset) by aligning the next token prediction to either $\mathcal{M}_{v1}$ or $\mathcal{M}_{v2}$. We define a masking (for individual tokens of a training sequence) following a simple heuristic depending on whether $\mathcal{M}^{C}_{v2}$ (the adapter being trained) predicts the correct tokens or not. If it does, we align to $\mathcal{M}_{v2}$ logits, otherwise, we align to $\mathcal{M}_{v1}$. The fine-tuning process of $\mathcal{M}^{C}_{v2}$ is depicted in Fig. 3. The fine-tuning loss to train the compatibility adapter, $\mathcal{L}_{comp}^{m}$, is defined below:
$m_{i}=\mathbf{1}[\text{argmax}~{}\sigma(z_{\mathcal{M}^{C}_{v2},i})\neq y_{i}]$ (1)
@@
Table 3: Compatible task adapter trained with MUSCLE (corresponding to metrics with suffix $c$) reduces negative flip rate. We see enhanced improvements for model updates that have smaller performance difference.
@@
Table 4: GSM8K math evaluation with exact match (EM). For compatibility adapter (corresponding to metrics with suffix $c$), we observe decreased negative flip rate while mostly maintaining performance gains.
@@ 5.1 Model Update Assumptions
To analyse the impact of model updates we consider parameter efficient fine-tuned models using Low-Rank Adapters (LoRA) Hu et al. (2021). Compared to previous work on continuous learning and model updates Qin et al. (2022, 2023), we do not limit model updates to be produced by only data updates, but consider different kinds of updates shown in Table 1. We include updates due to data, increased parameters, or different training strategies.
More importantly, we include wide range of downstream tasks to evaluate model compatibility, including generative tasks as summarized in Table 1. For all tasks we learn a LoRA adapter autoregressively (next-token prediction). Some of these tasks can be evaluated with classification-like approaches, e.g., by choosing the maximum log-likelihood of the possible answers (Gao et al., 2023), such as commonsense reasoning tasks (Zellers et al., 2019; Bisk et al., 2019). For other tasks such as dialogue summarization (Gliwa et al., 2019), we cannot evaluate in a classification-like manner.
@@ 5.2 Task Adapter Training
For each task and each old model $\mathcal{M}_{v1}$ and new model $\mathcal{M}_{v2}$, we train a LoRA adapter on all linear layers with $r=128$ and $\alpha=256$. We use a 0.8/0.2 training/validation split for 10 epochs. We choose the model by cross-entropy validation loss.
@@ 5.3 Compatibility Adapter Training
We keep all hyper-parameters the same to task adapter training, but train compatibility adapter with the $\mathcal{L}_{comp}^{m}$ loss defined in Eq. 2. We analyze statistical significance of the results on a subset of compatibility adapter training for Phi 1 to Phi 1.5 updates on PIQA using 3 random seeds. We observe a standard deviation of accuracy of 0.0012.
We compare $\mathcal{L}_{comp}^{m}$ with different masking strategies $m$ in the ablation studies in Section 6.5, but find that the version introduced in Section 4.1 works best for most model updates. Further, for model updates with large performance gaps between $\mathcal{M}_{v1}$ and $\mathcal{M}_{v2}$ we find that an auxiliary cross-entropy loss enhances stability of training for PIQA and SAMSum from Phi 1 to Phi 1.5, and all Phi updates in GSM8k and Hellaswag.
Figure 4: When updating LLM models (e.g. Llama 1 $\rightarrow$ Llama 2), we observe negative flips in different tasks. The smaller the performance gap from an old model to a new model, the more negative flips we observe. When evaluating continuous metrics with absolute ROUGE-1 value for summarization on SAMsum Gliwa et al. (2019), we observe a large fraction of negative flips. We show the exact models analysed in Table 8.
@@ 6.1 Negative Flips Occur in Model Updates
Fig. 4 shows that significant negative flips (up to more than 60%) exist for a variety of base model update scenarios and downstream tasks. We observe negative flips in updates within one model family (e.g. Llama/Vicuna, and Phi models). We find more negative flips for model updates with a smaller delta in performance gain. For generative tasks like SAMsum dialogue summarization, we observe a large number of negative flips, $\text{NFR}_{s}$, as continuous metrics are more sensitive to small changes when updated.
@@ 6.2 Reduced Negative Flips in Classification
Binary and multiple choice generative tasks can be evaluated with classification metrics, such as calculating the log likelihood of each answer option (we use Gao et al. (2023) for this use case). This can be interpreted as a “classical” sense of evaluating negative flips. In Table 3, we observe that MUSCLE decreases negative flips as shown when compared to regular model update without compatibility specific training ($\mathcal{M}_{v2}$). Specifically, we show reduction of NFR by 40% for the update of LLama 1 to Llama 2 and 39% for Vicuna 1.3 to Vicuna 1.5. For updates with large accuracy gap (for example Phi 1 to Phi 1.5), we observe a less strong enhancement with a negative flip rate reduction by 1-5%. In addition to reduction of negative flips we also observe an increased accuracy of up to 7% for Llama and Vicuna updates.
Similar to multiple-choice tasks, we can evaluate math questions in a classification-like approach to determine negative flips. In exact match (EM) evaluation we match the final result of the math question from the prediction to ground truth. Results are shown in Table 4. In this case, we observe that we can reduce the number of negative flips by 29% for Phi 1.5 to Phi 2. When the version 1 model is significantly less accurate (e.g., 3.4% for Phi 1), we observe a reduction in accuracy while only being able to decrease negative flips by 2%. For all other updates, MUSCLE increases expected-match accuracy while reducing negative flips.
Figure 5: Comparison of NFR vs NFRmc metrics to evaluate inconsistency when updating LLMs for HellaSwag task.
6.3 Increased Consistent Behavior
When we cannot achieve a positive flip (switching from an incorrect to a correct answer), we might prefer to at least maintain consistent behavior to old model to avoid unexpected behavior for the user. We evaluate negative flips and inconsistency flips. For model updates that have a large performance gap and a small number of negative flips to begin with, we see limited reduction in inconsistency flips. However, we observe that we can reduce the inconsistency flips for model updates with smaller accuracy gap such as the updates for Llama and Vicuna (Fig. 5) on the HellaSwag dataset.
@@ 6.3 Increased Consistent Behavior
Table 5: For summarization (SAMsum) generative task we reduce regression of ROUGE-1 score (R1) up to 27%.
6.4 Reduced Regression in Generative Tasks
When evaluating generative tasks, such as summarization, we observe that we can reduce regression of ROUGE-1 score performance (Table 5). We can reduce negative flips by 18% for updates with weaker version 1 models (Phi 1 $\rightarrow$ Phi 1.5) and 22% for smaller model updates (Vicuna 1.3 $\rightarrow$ Vicuna 1.5) and 27% for Phi 1.5 $\rightarrow$ Phi 2. When looking at average gain of ROUGE 1 Score we see that we decrease ROUGE-1 regression by 1-3%, while maintaining gain.
@@ 6.4 Reduced Regression in Generative Tasks
Table 6: A model update from Llama 1 $\rightarrow$ Llama 2 on PIQA. Different ablations and masking strategies and their impact on negative flips, positive flips and accuracy improvement.
6.5 The Effect of Different Masking Strategies
In this section, we analyze different training and masking strategies to evaluate our design choices. On the PIQA dataset and model update Llama 1 $\rightarrow$ Llama 2, we compare MUSCLE with different masking strategies. Intuitively, likelihood based masking strategies can be useful for tasks that are evaluated with log-likelihoods (e.g. PIQA, HellaSwag) where multiple choices are available. We compare likelihood based masking for example on individual tokens, $m_{i}=LL_{L}=\mathbf{1}[\sigma(z_{\mathcal{M}^{C}_{v2},i})<\sigma(z_{\mathcal{M}_{v1},i})]$ to check if the likelihood of the correct next token in the current model is smaller than the old model. Only in this case, we align to $\mathcal{M}_{v1}$, and align to $\mathcal{M}_{v2}$ for every other token. Alternatively, we can also compare the likelihood of the entire sequence $m_{i}=LL_{S}=\mathbf{1}[\sum_{i}[\sigma(z_{\mathcal{M}^{C}_{v2},i})<\sigma(z_{\mathcal{M}_{v1},i})]]$, such that we mask all tokens per sequence to get instance-wise masking. In both cases, we see a reduction in negative flips. However, likelihood based masking requires auxiliary cross-entropy loss for stability.
Furthermore, aligning to the old model only if it is already correct with a mask ($m_{i}=\mathbf{1}[\sigma(z_{\mathcal{M}_{v2},i}))=y_{i}]$) or purely (no masking) with KL ($a_{\mathcal{M}_{v1}}$) also lead to a small reduction in negative flips. Aligning to the old model if the current prediction is incorrect, i.e., Eq. 1, leads to the best reduction in negative flips while providing the biggest accuracy gain and increased positive flips (Table 6). For this best-performing strategy, we see that including cross entropy loss does not lead to performance gains. This strategy has the additional advantage that it takes into account extended inconsistent flips like explained in Fig. 5.
@@ 7 Conclusion
In this work, we study the task-specific compatibility problem when updating LLMs. We show that LLM updates with different scenarios, e.g., changes in model architecture, optimization, or training dataset, exhibit significant negative flips– instances previously classified or generated correctly, and incorrectly after the model update. We extend negative flip metric for discriminative and, for the first time, generative tasks, and report results for various models and tasks.
We propose a novel method (MUSCLE) to train task-specific compatibility adapter when updating an old LLM to a new LLM to reduce negative flips while maintaining performance gain. Our proposed method does not require a modification to the base model training, and is only based on adapter training. Further, as opposed to previous works, the proposed solution does not require both versions of the model in memory to enable compatibility, which is often infeasible due to large size of LLMs.
We observe a mitigation of negative flips of 40% for multiple-choice type evaluations, and 27% for continuous summarization evaluation. We also show insights into model properties that facilitate transfer, finding that our alignment masking strategy provides best results with an additional benefit of mitigating inconsistent update behavior.
8 Limitations
In the current study we do not consider a model update that includes changes in tokenization and/or vocabulary size (e.g.Llama 2 Touvron et al. (2023b) to LLama 3 AI@Meta (2024)). Future work can explore compatible vocabulary mapping strategies before learning from prior model versions. Our results indicate that the utility of aligning to prior model versions is limited by the performance of the prior model version. For example, see the update from Phi 1 to Phi 1.5 in Table 4, where Phi 1 only has an accuracy of 3%. In this case, it is arguable if an alignment to $\mathcal{M}_{v1}$ is desired and if the strive for compatibility outweighs a possible performance drop.
9 Ethical Considerations and Risks

### v2

@@ 4 Knowledge Transfer
Now that we have metrics to indicate model update regression, we propose a knowledge distillation approach to minimize this regression for the task-specific models $\mathcal{M}_{v1}$ and $\mathcal{M}_{v2}$. Typically, knowledge distillation minimizes the KL divergence between the soft targets $\sigma(z_{t})$ and $\sigma(z_{s})$, where $z_{s}$ and $z_{t}$ are logits predicted by student and teacher models, respectively.
@@
$,i$ denotes the $i$’th token, $n$ is the total number of tokens available for training, $T$ is temperature parameter, and $\sigma$ denotes softmax. Most knowledge distillation works consider the distillation from a trained teacher to an untrained student Tian et al. (2022); Rajasegaran et al. (2020). Recent work (Roth et al., 2024) tackles the goal of knowledge transfer between pre-trained student and teacher models while retaining student knowledge gained a priori, and shows that standard knowledge distillation between pre-trained models struggles to transfer knowledge without performance drops. Complementary to this work focusing on performance and maintaining prior knowledge, we tackle compatibility with prior models through knowledge transfer.
@@ 4.1 Model Update Strategy for Compatible LLM Evolution (MUSCLE)
When the base model is updated, we train a task-specific fine-tuned model, $\mathcal{M}^{C}_{v2}$, that has the accuracy benefits of $\mathcal{M}_{v2}$, but with most compatibility with $\mathcal{M}_{v1}$. We obtain $\mathcal{M}^{C}_{v2}$ by training a compatibility adapter applied to the base model $\mathcal{M}^{\text{base}}_{v2}$. We use knowledge from task-specific fine-tuned models $\mathcal{M}_{v1}$ and $\mathcal{M}_{v2}$ when training $\mathcal{M}^{C}_{v2}$. $\mathcal{M}_{v2}$ typically has increased prediction capabilities over $\mathcal{M}_{v1}$ (due to improvements in the base model), but $\mathcal{M}_{v1}$ has information on already correctly predicted tokens or instances that we want to minimize degradation towards.
We initialize the compatibility adapter with the task-specific adapter of $\mathcal{M}_{v2}$, and further fine-tune it (using the task training dataset) by aligning the next token prediction to either $\mathcal{M}_{v1}$ or $\mathcal{M}_{v2}$. We define masking (for individual tokens of a training sequence) following a simple heuristic depending on whether $\mathcal{M}^{C}_{v2}$ (the adapter being trained) predicts the correct tokens or not. If it does, we align to $\mathcal{M}_{v2}$ logits, otherwise, we align to $\mathcal{M}_{v1}$. The fine-tuning process of $\mathcal{M}^{C}_{v2}$ is depicted in Fig. 3. The fine-tuning loss to train the compatibility adapter, $\mathcal{L}_{comp}^{m}$, is defined below:
$m_{i}=\mathbf{1}[\text{argmax}~{}\sigma(z_{\mathcal{M}^{C}_{v2},i})\neq y_{i}]$
@@
Table 4: Compatible task adapter trained with MUSCLE (corresponding to metrics with suffix $c$) reduces negative flip rate on the test sets of multiple-choice language tasks. We see most improvements for model updates that have small performance differences.
@@
Table 5: GSM8K math evaluation with exact match (EM) over the test dataset. For compatibility adapter (corresponding to metrics with suffix $c$), we observe a decreased negative flip rate while mostly maintaining performance gains.
@@ 5.1 Model Update Assumptions
To analyze the impact of model updates we consider parameter-efficient fine-tuned models using Low-Rank Adapters (LoRA) Hu et al. (2021). Compared to previous work on continuous learning and model updates Qin et al. (2022, 2023), we do not limit model updates to be produced by only data updates, but consider different kinds of updates shown in Table 2. We include updates due to data, increased parameters, or different training strategies. We include a wide range of downstream tasks to evaluate model compatibility, including generative tasks as summarized in Table 2. For all tasks, we learn the LoRA adapter autoregressively (next-token prediction).
@@ 5.2 Task Adapter Training
For each task and each old model $\mathcal{M}_{v1}$ and new model $\mathcal{M}_{v2}$, we train a LoRA adapter on all linear layers with $r=128$ and $\alpha=256$. We use a 0.8/0.2 training/validation split for 10 epochs. We choose the model by cross-entropy validation loss. More information on hyperparameters is shown in Table 8.
@@ 5.3 Compatibility Adapter Training
We keep all hyper-parameters the same for task adapter training, and train compatibility adapter with the $\mathcal{L}_{comp}^{m}$ loss defined in Eq. 1. We analyze the statistical significance of the results on a subset of compatibility adapter training for Phi 1 to Phi 1.5 updates on PIQA using 3 random seeds. We observe a standard deviation of accuracy of 0.0012.
We compare $\mathcal{L}_{comp}^{m}$ with different masking strategies $m$ in the ablation studies in Section 6.5, but find that the version introduced in Section 4.1 works best for most model updates. For model updates with large performance gaps between $\mathcal{M}_{v1}$ and $\mathcal{M}_{v2}$ we find that an auxiliary cross-entropy loss enhances the stability of training for PIQA and SAMSum from Phi 1 to Phi 1.5, and all Phi updates in GSM8k and Hellaswag (more in Section A.4).
5.4 Similarity Metrics
We evaluate multiple-choice tasks with classification-like approaches by choosing the maximum log-likelihood of the possible answers (Gao et al., 2023). For math tasks, we use exact-match accuracy of the final calculated result. For summarization, we cannot evaluate in a classification-like manner. We use ROUGE-1  (Lin, 2004), given that we do not observe relative ranking differences for different ROUGE-n.
@@ 6.1 Negative Flips Occur in Model Updates
Fig. 4 shows that significant negative flips (up to more than 60%) exist for a variety of base model update scenarios and downstream tasks. We observe negative flips in updates within one model family (e.g. Llama/Vicuna, and Phi models). We find more negative flips for model updates with a smaller delta in performance gain. For generative tasks like SAMsum dialogue summarization, we observe a large number of negative flips as continuous metrics are more sensitive to small changes when updated.
@@ 6.2 Reduced Negative Flips in Classification
Figure 4: When updating LLM models (e.g. Llama 1 $\rightarrow$ Llama 2), we observe negative flips in different tasks. The smaller the performance gap from an old model to a new model, the more negative flips we observe. We indicate the performance gap by the difference in exact match for GSM8K, Rouge-1 for SAMSum, and log-likelihood-based accuracy for PIQA and HellaSwag. When evaluating continuous metrics with absolute ROUGE-1 value for summarization on SAMSum, we observe a large fraction of negative flips. We show the exact models analyzed in Table 9.
In Table 4, we observe that MUSCLE decreases negative flips when compared to regular model updates without compatibility-specific training ($\mathcal{M}_{v2}$). Specifically, we see a reduction of NFR by 40% for the update of LLama 1 to Llama 2 and 39% for Vicuna 1.3 to Vicuna 1.5. For updates with a large performance gap (for example Phi 1 to Phi 1.5), we observe a less strong enhancement with a negative flip rate reduction by 1-5%. In addition to the reduction of negative flips, we also observe an increased accuracy of up to 7% for Llama and Vicuna updates.
In exact match (EM) evaluation, we match the final result of the math question from the prediction to ground truth. Results are shown in Table 5. In this case, we observe that we can reduce the number of negative flips by 29% for Phi 1.5 to Phi 2. When the version 1 model is significantly less accurate (e.g., 3.4% exact-match accuracy for Phi 1), we observe a reduction in accuracy with the compatibility adapter while only being able to decrease negative flips by 2%. For all other updates, MUSCLE increases expected-match accuracy while reducing negative flips.
@@
Table 6: For summarization (SAMsum) generative task we reduce model update regression of ROUGE-1 score (R1) by up to 27%.
6.3 Increased Consistent Behavior
When we cannot achieve a positive flip (switching from an incorrect to a correct answer), we might prefer to at least maintain consistent behavior to the old model to avoid unexpected behavior for the user. We evaluate negative flips rate (NFR) and inconsistency flips rate ($\text{NFR}_{mc}$). For model updates that have a large performance gap and a small number of negative flips to begin with (Phi models), we see a limited reduction in inconsistency flips. However, we observe that we can reduce the inconsistency flips with our method for model updates with small accuracy gaps such as the updates for Llama and Vicuna (Fig. 5) on the HellaSwag dataset.
Figure 5: Comparison of NFR vs NFRmc metrics to evaluate inconsistency when updating LLMs for HellaSwag task. We see that using our compatibility adapter (denoted by c), we can reduce inconsistency for Llama and Vicuna models.
6.4 Reduced Model Update Regression in Generative Tasks
When evaluating generative tasks, we can reduce model update regression of ROUGE-1 score performance (Table 6). We can reduce negative flips by 18% for updates with weaker version 1 models (Phi 1 $\rightarrow$ Phi 1.5) and 22% for smaller model updates (Vicuna 1.3 $\rightarrow$ Vicuna 1.5) and 27% for Phi 1.5 $\rightarrow$ Phi 2. We see that we decrease ROUGE-1 model update regression by 1-3%, while maintaining gain.
6.5 The Effect of Different Masking Strategies
We analyze different training and masking strategies to evaluate our design choices. On the PIQA dataset and model update Llama 1 $\rightarrow$ Llama 2, we compare MUSCLE with different masking strategies. Intuitively, instance or token-wise likelihood-based masking strategies can be useful for tasks that are evaluated with log-likelihoods (e.g. PIQA, HellaSwag) where multiple choices are available. We compare likelihood-based masking for example on individual tokens,
$m_{i}=LL_{L}=\mathbf{1}[\sigma(z_{\mathcal{M}^{C}_{v2},i})<\sigma(z_{\mathcal{M}_{v1},i})]$ (2)
to check if the likelihood of the ground-truth next token in the current model is smaller than the old model. Only in this case, we align to $\mathcal{M}_{v1}$, and align to $\mathcal{M}_{v2}$ for every other token. Alternatively, we can also compare the likelihood of the entire sequence
$m_{i}=LL_{S}=\mathbf{1}[\sum_{i}[\sigma(z_{\mathcal{M}^{C}_{v2},i})<\sigma(z_{\mathcal{M}_{v1},i})]]$ (3)
such that we mask all tokens per sequence to get instance-wise masking. For both of these strategies, we see a reduction in negative flips. We note that likelihood-based masking requires auxiliary cross-entropy loss for stability (A.4).
@@ 6.5 The Effect of Different Masking Strategies
Table 7: A model update from Llama 1 $\rightarrow$ Llama 2 on PIQA. Different ablations and masking strategies and their impact on negative flips, positive flips, and accuracy improvement.
Aligning to the old model only if it is already correct with a mask $\mathcal{L}_{comp}^{\mathcal{M}_{v1}=y}$ with $m_{i}=\mathbf{1}[\sigma(z_{\mathcal{M}_{v1},i})=y_{i}]$ or without masking with KL ($a_{\mathcal{M}_{v1}}$) leads to a small reduction in negative flips. Our approach that aligns to the old model if the current prediction is incorrect ($\mathcal{L}_{comp}^{\mathcal{M}^{C}_{v2}\neq y}$), leads to the best reduction in negative flips while providing the biggest accuracy gain and increased positive flips (Table 7). This strategy has the additional advantage that it takes into account extended inconsistent flips explained in Fig. 5, as it aligns to $\mathcal{M}_{v1}$ whenever $\mathcal{M}_{v2}$ is incorrect. For this best-performing strategy, we see that including cross-entropy loss (CE+$\mathcal{L}_{comp}^{\mathcal{M}^{C}_{v2}\neq y}$) does not lead to additional performance gains.
6.6 Behavior for Different Model Pairs
A general point in the analysis of negative flips is its connection with accuracy: if accuracy is 100%, the NFR must be 0%. This means that compatible model updates become harder when updating $\mathcal{M}_{v1}\rightarrow\mathcal{M}_{v2}$ with lower $\mathcal{M}_{v2}$ accuracy. We report observed compatibility when using vanilla adapter learning, and show a reduction in NFR when using the proposed method (MUSCLE). When $\text{acc}(\mathcal{M}_{v2})>\text{acc}(\mathcal{M}_{v1})$ for a task, the observed NFR is small (Fig. 4). This has been the case irrespective of the model update scenario and task. We find that MUSCLE reduces the observed NFR the most when $\text{acc}(\mathcal{M}_{v2})\approx\text{acc}(\mathcal{M}_{v1})$. In such cases, the incompatibility between vanilla model updates is not dominated by the fact that one model version is more accurate than the other, making it an easier use case to be improved by our proposed distillation loss. With multiple frequent incremental model updates in academia and industry, this is a very relevant use case to focus on.
@@ 7 Conclusion
In this work, we study the task-specific compatibility problem when updating LLMs. We show that LLM updates with different scenarios, e.g., changes in model architecture, optimization, or training dataset, exhibit significant negative flips – instances previously classified or generated correctly, and incorrectly after the model update. We extend the negative flip metric for discriminative and, for the first time, generative tasks, and report results for various models and tasks.
We propose a novel method (MUSCLE) to train task-specific compatibility adapters when updating an old LLM to a new LLM to reduce negative flips while maintaining performance gain. Our proposed method does not require a modification to the base model training and is only based on adapter training. Further, as opposed to previous works, the proposed solution does not require both versions of the model in memory to enable compatibility, which is often infeasible due to the large size of LLMs.
We observe a mitigation of negative flips of 40% for multiple-choice type evaluations, and 27% for continuous summarization evaluation. We also show insights into model properties that facilitate transfer, finding that our alignment masking strategy provides best results with the additional benefit of mitigating inconsistent update behavior.
8 Limitations, Risks and Future Work
We do not consider a model update that includes changes in tokenization and/or vocabulary size (e.g.Llama 2 Touvron et al. (2023b) to LLama 3 AI@Meta (2024)). Future work can explore compatible vocabulary mapping strategies before learning from prior model versions.
Tackling Large Performance Gaps
When there is a large performance gap between $M_{v1}$ and $M_{v2}$, loss hyperparameter weighting could be an interesting avenue to explore. We unsuccessfully experimented with simple weighting based on average accuracy. We hypothesize that instance-based loss weighting could be a more promising approach to tackle this model update case. As this would require intensive experimentation, we leave this as an avenue for future work. In general, the utility of aligning to prior model versions is limited by the performance of the prior model version. For example, see the update from Phi 1 to Phi 1.5 in Table 5, where Phi 1 only has an accuracy of 3%. In this case, it is arguable if an alignment to $\mathcal{M}_{v1}$ is desired and if the strive for compatibility outweighs a possible performance drop.
MUSCLE Performance Improvements
In our work, we observe an interesting overall performance improvement following knowledge transfer, which is an intriguing finding regarding the possibilities of this line of work. Previous works on model compatibility using distillation-like losses have also observed such a phenomenon (e.g., Jaeckle et al. (2023)). Given that we initialize the compatibility adapter with $M_{v2}$ (an independently trained task adapter) and continue fine-tuning it with knowledge transfer loss from $M_{v1}$, one can argue that the observed performance improvement demonstrates an ensemble knowledge effect (i.e., knowledge from both $M_{v2}$ and $M_{v1}$ is aggregated into our compatibility adapter).
Ethical Considerations and Risks